# A. Abed Abud CERN, The European Organization for Nuclear Research, 1211 Meyrin, Switzerland B. Abi

### v1

@@ Abstract
The determination of the direction of a stellar core collapse via its neutrino emission is crucial for the identification of the progenitor for a multimessenger follow-up. A highly effective method of reconstructing supernova directions within the Deep Underground Neutrino Experiment (DUNE) is introduced. The supernova neutrino pointing resolution is studied by simulating and reconstructing electron-neutrino charged-current absorption on and elastic scattering of neutrinos on electrons. Procedures to reconstruct individual interactions, including a newly developed technique called “brems flipping”, as well as the burst direction from an ensemble of interactions are described.
Performance of the burst direction reconstruction is evaluated for supernovae happening at a distance of 10 kpc for a specific supernova burst flux model.
The pointing resolution is found to be 3.4 degrees at 68% coverage for a perfect interaction-channel classification and a fiducial mass of 40 kton, and 6.6 degrees for a 10 kton fiducial mass respectively. Assuming a 4% rate of charged-current interactions being misidentified as elastic scattering, DUNE’s burst pointing resolution is found to be 4.3 degrees (8.7 degrees) at 68% coverage.
@@
Detecting neutrinos from core-collapse supernovae opens up the possibility to study the astrophysics of stellar collapse and the properties of neutrinos and their interactions. The first and only supernova neutrino detection on Earth so far was achieved with SN1987A, when a few tens of electron-antineutrino events111We note here the use of “event” to refer to an individual recorded neutrino interaction, as per standard particle physics usage. In this paper, “burst” will refer to the ensemble of events from a single core collapse. were registered [1, 2, 3]. While it was already possible to derive valuable insights from these data (e.g., Ref. [4, 5]), a high-statistics, high-resolution detection of various neutrino flavors from a future supernova will bring invaluable advancement in particle physics and astrophysics.
Information can be derived from the detected neutrino event rate, timing, energy spectrum, flavor composition, and angular distribution.
@@
An important characteristic feature of a supernova neutrino burst is that it emerges promptly after the core bounce, preceding the related electromagnetic phenomena. As neutrinos interact only via the weak interaction, they can escape more easily in comparison to photons, allowing the neutrino signal to be observed well before the associated electromagnetic radiation (as can gravitational radiation). The Supernova Early Warning System (SNEWS) [6] is designed such that information retrieved from the supernova neutrino signal can be made available worldwide quickly, thereby facilitating the prompt detection of subsequent multi-messenger supernova-related phenomena [7, 8].
@@
Clearly, it is highly desirable for information about the position of the supernova in the sky to be ascertained as fast as possible. Not only is the direction useful for the localization of the supernova to enable prompt measurements, but for the case when the core collapse fails to produce a bright explosion in electromagnetic radiation (e.g. black hole formation), a neutrino burst direction may help to locate a missing progenitor using archival data.
@@
Such pointing information is available from the neutrinos themselves. One possible strategy for determining the direction to the supernova is via “triangulation” from the relative timing of neutrinos observed at different locations around the globe [9, 10, 11, 12]. However, the most promising way to achieve precision pointing is to exploit anisotropic neutrino interactions [9] for which information about the incoming neutrino direction is preserved in the final-state particle angular distribution. Water Cherenkov detectors such as Super-Kamiokande and Hyper-Kamiokande can make use of directional Cherenkov radiation to determine the supernova direction [13].
@@
The Deep Underground Neutrino Experiment (DUNE) is capable of a supernova burst detection among other physics goals [14, 15]. An overview of DUNE’s supernova detection capability is given in Ref. [14].
In the several tens of kilotonnes (kton) of liquid argon (LAr) volume of DUNE, charged-current interactions of electron neutrinos (CC) on 40Ar and neutrinos of all flavors scattering elastically on electrons (eES) will result in charge signals in DUNE’s time projection chamber (TPC) and scintillation light that is read out via photosensors.
@@
In this work, both CC and eES interactions are discussed.222We focus here on directional information derived from the expected prompt burst of tens-of-MeV neutrinos, although we note that higher energy (GeV scale or more) neutrinos may follow a supernova [16, 17]. These may provide precision pointing due to both higher intrinsic lepton-neutrino collimation and better detector performance at high energy; however, they likely arrive on a long time scale (hours to years) after the supernova. Primarily eES interactions carry directional information, but the near-isotropic CC interactions have a higher cross section. A supernova burst alert from DUNE, including pointing information, will be a valuable input to SNEWS.
@@
The overall ability for DUNE to point to a supernova using the events recorded from a core collapse depends on several factors. Two primary factors affect the direction resolution of individually recorded neutrino events. First, there are intrinsic angular spreads between the recoil directions of the final-state products and the neutrino direction. For eES interactions, which are the most important for the pointing ability, this energy-dependent spread is very well understood from weak interaction physics; for CC it is less well known. Second, the detector angular resolution smears the reconstructed direction with respect to the final-state electron direction. Detector-resolution smearing can in principle be improved with better reconstruction algorithms, although there will be an intrinsic physical limit for a given detector configuration. For this study, we use standard DUNE reconstruction techniques (detailed in Sec. IV), with some minor improvements, as well as a novel additional technique we call “brems flipping,” which provides event-by-event a modest improvement in head-tail directional disambiguation by looking at the location of the secondary particles’ tracks relative to the primary track.
@@
While individual-event resolution on neutrino direction is relatively modest due to both physical angular spread and instrumental uncertainty, for the statistical ensemble of events in a burst the directional determination improves approximately by the inverse square root of the number of events detected. We evaluate here the overall supernova burst pointing resolution, which will enable a successful contribution from DUNE to a multi-messenger detection of a core-collapse supernova. The scope of this study, which makes use of standard offline reconstruction software, addresses only intrinsic pointing ability and does not consider latency for dissemination of pointing information.
@@ II Supernova neutrino emission
When a massive star has depleted all of its nuclear fuel, the outgoing radiation pressure ceases to counteract the inward gravitational pull of the star, collapsing it into a compact object such as a neutron star or a black hole. During this process, 99% of the gravitational binding energy of the remnant is emitted in the form of neutrinos with energies of a few tens of MeV over a few tens of seconds. Supernova neutrinos are released in several stages during a core collapse. At the beginning of the collapse (over tens of milliseconds), neutrinos are primarily produced via electron capture () as the star undergoes neutronization. During the subsequent accretion phase (which lasts for tens to hundreds of milliseconds), more neutrinos of all flavors are created, counteracting the shock heating of the in-falling matter. Subsequently, pairs are emitted over the next tens of seconds, such that these neutrino pairs shed most of the gravitational binding energy, thereby cooling the remnant [18].
@@
The study of supernova neutrinos is particularly useful in providing insights on varied topics in astrophysics and neutrino physics. As neutrinos are intimately involved in the collapse and subsequent explosion processes, measurements of the supernova neutrino signal allow for the examination of the complex interactions that occur during the core collapse. Additionally, supernova neutrino signals have the important feature that the initial luminosity is roughly equally divided among flavors. The subsequent flavor transitions provide information on the neutrino mass ordering, as well as insights into exotic flavor transition physics [19, 20, 21, 22].
@@
For this study, the GKVM [23] core-collapse supernova neutrino emission model is used to describe the neutrino energy spectrum. There are known to be fairly large (up to an order of magnitude) uncertainties on the supernova neutrino flux prediction, and result both from intrinsic progenitor variance and from theoretical uncertainties. These variances will have an impact on the total number of neutrinos detected as well as on details of the neutrino flavor composition and spectra. However, the pointing capabilities of the DUNE detector are primarily sensitive to event statistics rather than to details of the model flux. We therefore do not survey different supernova models here. The selected model produces an intermediate number of eES events among a range of models. Expected event rates without energy smearing are calculated with SNOwGLoBES [24], which computes the recoil energy distribution as described in the next section. It is assumed that the supernova explosion occurs at a distance of 10 kpc from Earth, within the Milky Way for our selected model. The total expected number of events and the events per interaction channel are presented in Tab. 1. We do not study the effects of flavor transitions333Flavor transitions may affect the total number of neutrino events, but with uncertainty not exceeding the overall flux model uncertainties; spectral modulations will be a subdominant effect on pointing capabilities..
@@ III Supernova neutrino detection at DUNE
DUNE is unique in the sense that it will be able to register large numbers of electron neutrinos, while all other experiments of similar size, e.g., Hyper-K [25] and JUNO [26], primarily detect electron antineutrinos via inverse beta decay on free protons. Immediately following the core collapse, electron-neutrino emission through neutronization dominates and insights on the neutrino mass ordering can be gained from observed differences in the neutrino flavor composition and spectra due to oscillation dynamics within the supernova [22].
@@
We consider neutrino detection at DUNE’s far detector liquid argon time projection chamber (LArTPC). The full planned design consists of 40 kton of fiducial mass444We note this is the fiducial mass for long-baseline physics; active mass for supernova event detection could potentially be larger. of liquid argon placed into four modules. Both single-module (10 kton) and all-modules (40 kton) performance is considered in this study. We consider DUNE’s first-module design for this study. Each of the active volumes is bound by the cathode plane assembly (CPA) and the anode plane assembly (APA) and surrounded by the field cage. A uniform electric field is created between the CPA and APA, drifting ionization charges toward the APA. Three planes of sensing wires, each with a different orientation, are located at the APA. Charge depositions on these readout planes are used to reconstruct the location of the particle energy deposition in two dimensions, and the drift timing information allows for the reconstruction of the third spatial coordinate. In addition to the LArTPCs, the Photon Detection System (PDS) is used to tag neutrino events. In this study, information on the timing of the interaction is retrieved from the detection of photon flashes to be able to estimate the charge loss during drift. More details about the photon detectors can be found in [27].
The two neutrino-interaction channels considered in this study are neutrino-electron elastic scattering interactions and CC absorption interactions.
@@
Neutral-current and anti-neutrino CC interactions are also expected [14]. However, as these channels are known to be subdominant and are at present not well understood, they are not included in this study. In the following, the nature of the dominant interaction channels is described and the simulated energy spectra and angular distributions in the detector are discussed. For eES events, the event rates are determined from SNOwGLoBES, using the GKVM supernova model. It also provides the recoil energy distribution as input to a LArSoft [28] neutrino-electron scattering event generator. For CC events, MARLEY [29] and its LArSoft interface are applied to generate the events with the correct energy distribution. Once again the GKVM supernova model is provided as input to MARLEY in this case. All simulations are done in a subset of the full DUNE far detector geometry simulating 1.6 kton of fiducial volume to reduce memory, disk, and computing time requirements.
The simulated workspace features six planes of APAs (two APAs tall) along the beam direction. For the full detector module, there are 150 APAs proposed in total, stacked two-tall and 25 along the beam direction. Because the spatial extent of the neutrino events under consideration is much smaller than the simulated workspace size, we expect the scaling down of the simulated geometry to have a negligible effect on the conclusions of the study.
Table 1: The eES and CC event rates (total for the interaction channel in bold) for DUNE within a fiducial volume of 40 kton, and for a core collapse at a distance of 10 kpc, using the GKVM model. No triggering efficiency or detector resolution effects are applied here, in contrast to [14]. Neutrino flavor transition effects beyond those included in the GKVM model are neglected. The neutrino event rate scales linearly with detector mass and with inverse-square dependence on supernova distance.
Channel
Expected Event Count
(all flavors)
325.8
155.5
67.3
55.2
47.8
3300.0
@@ III.1 Neutrino-Electron Elastic Scattering
The most relevant interaction for directional information is when a neutrino elastically scatters off of an electron in the LAr. This interaction applies to all flavors but the largest cross section is for . A visualization of a Geant4 event is shown in Fig. 1.
Figure 1: Geant4 illustration of the energy deposition for examples of two event types: (a) CC: A 25 MeV electron neutrino is absorbed by an argon nucleus resulting in the excitation of the nucleus and the emission of an electron. (b) eES: An incoming electron neutrino of 12 MeV scatters elastically in the LAr. Charged particles are emitted in the forward direction of the primary electron. The CC primary electron tracks are on average longer than eES tracks due to the lower energies of the eES recoils; in contrast, CC electrons tend to retain most of the energy of the incoming neutrino.
Gamma tracks are not shown in the display, but red blips (representing electrons) from gamma interactions with the argon (primarily Compton scatters) can be seen.
The direction of the scattered electron is highly correlated to the direction of the neutrino. In particular, the scattering angle is described by [30]:
where is the electron kinetic energy, is the neutrino energy, and is the electron mass.
The distribution of is given by the differential cross section:
where is the Fermi coupling constant, and and are given according to neutrino flavor [30] in Tab. 2.
Table 2: Coupling strengths in the cross section of neutrino-electron scattering from Ref. [30]. The cross section of electron scattering is enhanced due to the possibility of both neutral- and charged-current interactions occurring; -electron elastic scattering is helicity-suppressed with respect to scattering.
Flavor

### v2

# The DUNE Collaboration Supernova Pointing Capabilities of DUNE

@@ Abstract
The determination of the direction of a stellar core collapse via its neutrino emission is crucial for the identification of the progenitor for a multimessenger follow-up. A highly effective method of reconstructing supernova directions within the Deep Underground Neutrino Experiment (DUNE) is introduced. The supernova neutrino pointing resolution is studied by simulating and reconstructing electron-neutrino charged-current absorption on and elastic scattering of neutrinos on electrons. Procedures to reconstruct individual interactions, including a newly developed technique called “brems flipping”, as well as the burst direction from an ensemble of interactions are described. Performance of the burst direction reconstruction is evaluated for supernovae happening at a distance of 10 kpc for a specific supernova burst flux model. The pointing resolution is found to be 3.4 degrees at 68% coverage for a perfect interaction-channel classification and a fiducial mass of 40 kton, and 6.6 degrees for a 10 kton fiducial mass respectively. Assuming a 4% rate of charged-current interactions being misidentified as elastic scattering, DUNE’s burst pointing resolution is found to be 4.3 degrees (8.7 degrees) at 68% coverage.
Contents
I Introduction
@@
Detecting neutrinos from core-collapse supernovae opens up the possibility to study the astrophysics of stellar collapse and the properties of neutrinos and their interactions. The first and only supernova neutrino detection on Earth so far was achieved with SN1987A, when a few tens of electron-antineutrino events111We note here the use of “event” to refer to an individual recorded neutrino interaction, as per standard particle physics usage. In this paper, “burst” will refer to the ensemble of events from a single core collapse. were registered [Bionta:1987qt, Hirata:1987hu, Alekseev:1987ej]. While it was already possible to derive valuable insights from these data (e.g., Ref. [burrows1987neutrinos, schramm1990new]), a high-statistics, high-resolution detection of various neutrino flavors from a future supernova will bring invaluable advancement in particle physics and astrophysics. Information can be derived from the detected neutrino event rate, timing, energy spectrum, flavor composition, and angular distribution.
@@
An important characteristic feature of a supernova neutrino burst is that it emerges promptly after the core bounce, preceding the related electromagnetic phenomena. As neutrinos interact only via the weak interaction, they can escape more easily in comparison to photons, allowing the neutrino signal to be observed well before the associated electromagnetic radiation (as can gravitational radiation.) The Supernova Early Warning System (SNEWS) [al2021snews] is designed such that information retrieved from the supernova neutrino signal can be made available worldwide quickly, thereby facilitating the prompt detection of subsequent multi-messenger supernova-related phenomena [Adams:2013ana, Nakamura:2016kkl].
@@
Clearly, it is highly desirable for information about the position of the supernova in the sky to be ascertained as fast as possible. Not only is the direction useful for the localization of the supernova to enable prompt measurements, but for the case when the core collapse fails to produce a bright explosion in electromagnetic radiation (e.g., black hole formation), a neutrino burst direction may help to locate a missing progenitor using archival data.
@@
Such pointing information is available from the neutrinos themselves. One possible strategy for determining the direction to the supernova is via “triangulation” from the relative timing of neutrinos observed at different locations around the globe [Beacom:1998fj, Muhlbeier:2013gwa, Brdar:2018zds, Linzer:2019swe]. However, the most promising way to achieve precision pointing is to exploit anisotropic neutrino interactions [Beacom:1998fj] for which information about the incoming neutrino direction is preserved in the final-state particle angular distribution. Water Cherenkov detectors such as Super-Kamiokande and Hyper-Kamiokande can make use of directional Cherenkov radiation to determine the supernova direction [abe2016real].
@@
The Deep Underground Neutrino Experiment (DUNE) is capable of a supernova burst detection among other physics goals [abi2021supernova, abi2020deep]. An overview of DUNE’s supernova detection capability is given in Ref. [abi2021supernova]. In the several tens of kilotonnes (kton) of liquid argon (LAr) volume of DUNE, charged-current interactions of electron neutrinos (CC) on 40Ar and neutrinos of all flavors scattering elastically on electrons (eES) will result in charge signals in DUNE’s time projection chamber (TPC) and scintillation light that is read out via photosensors. In this work, both CC and eES interactions are discussed.222We focus here on directional information derived from the expected prompt burst of tens-of-MeV neutrinos, although we note that higher energy (GeV scale or more) neutrinos may follow a supernova [Tomas:2003xn, Murase:2017pfe]. These may provide precision pointing due to both higher intrinsic lepton-neutrino collimation and better detector performance at high energy; however, they likely arrive on a long time scale (hours to years) after the supernova. Primarily eES interactions carry directional information, but the near-isotropic CC interactions have a higher cross section. A supernova burst alert from DUNE, including pointing information, will be a valuable input to SNEWS.
@@
The overall ability for DUNE to point to a supernova using the events recorded from a core collapse depends on several factors. Two primary factors affect the direction resolution of individually recorded neutrino events. First, there are intrinsic angular spreads between the recoil directions of the final-state products and the neutrino direction. For eES interactions, which are the most important for the pointing ability, this energy-dependent spread is very well understood from weak interaction physics; for CC it is less well known. Second, the detector angular resolution smears the reconstructed direction with respect to the final-state electron direction. Detector-resolution smearing can in principle be improved with better reconstruction algorithms, although there will be an intrinsic physical limit for a given detector configuration. For this study, we use standard DUNE reconstruction techniques (detailed in Sec. IV), with some minor improvements, as well as a novel additional technique we call “brems flipping”, which provides event-by-event a modest improvement in head-tail directional disambiguation by looking at the location of the secondary particles’ tracks relative to the primary track.
@@
While individual-event resolution on neutrino direction is relatively modest due to both physical angular spread and instrumental uncertainty, for the statistical ensemble of events in a burst the directional determination improves approximately by the inverse square root of the number of events detected. We evaluate here the overall supernova burst pointing resolution, which will enable a meaningful contribution from DUNE to a multi-messenger detection of a core-collapse supernova. The scope of this study, which makes use of standard offline reconstruction software, addresses only intrinsic pointing ability and does not consider latency for dissemination of pointing information.
@@ II Supernova neutrino emission
When a massive star has depleted all of its nuclear fuel, the outgoing radiation pressure ceases to counteract the inward gravitational pull of the star, collapsing it into a compact object such as a neutron star or a black hole. During this process, 99% of the gravitational binding energy of the remnant is emitted in the form of neutrinos with energies of a few tens of MeV over a few tens of seconds. Supernova neutrinos are released in several stages during a core collapse. At the beginning of the collapse (over tens of milliseconds), neutrinos are primarily produced via electron capture () as the star undergoes neutronization. During the subsequent accretion phase (which lasts for tens to hundreds of milliseconds), more neutrinos of all flavors are created, counteracting the shock heating of the in-falling matter. Subsequently, pairs are emitted over the next tens of seconds, such that these neutrino pairs shed most of the gravitational binding energy, thereby cooling the remnant [scholberg2012supernova].
@@
The study of supernova neutrinos is particularly useful in providing insights on varied topics in astrophysics and neutrino physics. As neutrinos are intimately involved in the collapse and subsequent explosion processes, measurements of the supernova neutrino signal allow for the examination of the complex interactions that occur during the core collapse. Additionally, supernova neutrino signals have the important feature that the initial luminosity is roughly equally divided among flavors. The subsequent flavor transitions provide information on the neutrino mass ordering, as well as insights into exotic flavor transition physics [duan2010collective, mirizzi2016supernova, gil2003oscillation, Scholberg:2017czd].
@@
For this study, the GKVM [gava2009dynamical] core-collapse supernova neutrino emission model is used to describe the neutrino energy spectrum. There are known to be fairly large (a factor of several) uncertainties on the supernova neutrino event rate prediction. These uncertainties result both from intrinsic progenitor variance and from theoretical uncertainties. These variances will have an impact on the total number of neutrinos detected as well as on details of the neutrino flavor composition and spectra. However, the pointing capabilities of the DUNE detector are primarily sensitive to event statistics rather than to details of the model flux. We therefore do not survey different supernova models here. The selected model produces an intermediate number of eES events among a range of models. Expected event rates without energy smearing are calculated with SNOwGLoBES [snowglobes], which computes the recoil energy distribution as described in the next section. It is assumed that the supernova explosion occurs at a distance of 10 kpc from Earth, within the Milky Way for our selected model. The total expected number of events and the events per interaction channel are presented in Tab. 1. We do not study the effects of flavor transitions333Flavor transitions may affect the total number of neutrino events, but with uncertainty not exceeding the overall flux model uncertainties; spectral modulations will be a subdominant effect on pointing capabilities..
@@ III Supernova neutrino detection at DUNE
DUNE is unique in the sense that it will be able to register large numbers of electron neutrinos, while all other experiments of similar size, e.g., Hyper-K [abe2018hyper] and JUNO [an2016neutrino], primarily detect electron antineutrinos via inverse beta decay on free protons. Immediately following the core collapse, electron-neutrino emission through neutronization dominates and insights on the neutrino mass ordering can be gained from observed differences in the neutrino flavor composition and spectra due to oscillation dynamics within the supernova [Scholberg:2017czd].
@@
We consider neutrino detection at DUNE’s far detector liquid argon time projection chamber (LArTPC) [abi2020volume]. The full planned design consists of 40 kton of fiducial mass444We note this is the fiducial mass for long-baseline physics; active mass for supernova event detection could potentially be larger. of liquid argon placed into four modules. Both single-module (10 kton) and all-modules (40 kton) performance is considered in this study. We consider the horizontal-drift module design for this study. Each of the active volumes is bound by the cathode plane assembly (CPA) and the anode plane assembly (APA) and surrounded by the field cage. A uniform electric field is created between the CPA and APA, drifting ionization charges toward the APA. Three planes of sensing wires, each with a different orientation, are located at the APA. Charge depositions on these readout planes are used to reconstruct the location of the particle energy deposition in two dimensions, and the drift timing information allows for the reconstruction of the third spatial coordinate. In addition to the LArTPCs, the Photon Detection System (PDS) is used to tag neutrino events. In this study, information on the timing of the interaction is retrieved from the detection of photon flashes to be able to estimate the charge loss during drift. More details about the photon detectors can be found in [Falcone:20220h].
@@
The two neutrino-interaction channels considered in this study are neutrino-electron elastic scattering interactions and CC absorption interactions. Neutral-current and anti-neutrino CC interactions are also expected [abi2021supernova]. However, as these channels are known to be subdominant and are at present not well understood, they are not included in this study. In the following, the nature of the dominant interaction channels is described and the simulated energy spectra and angular distributions in the detector are discussed. For eES events, the event rates are determined from SNOwGLoBES, using the GKVM supernova model. It also provides the recoil energy distribution as input to a LArSoft [snider2017larsoft] neutrino-electron scattering event generator. For CC events, MARLEY [gardiner2021simulating] and its LArSoft interface are applied to generate the events with the correct energy distribution. Once again the GKVM supernova model is provided as input to MARLEY in this case. All simulations are done in a subset of the full DUNE far detector geometry simulating 1.6 kton of fiducial volume to reduce memory, disk, and computing time requirements. The simulated workspace features six planes of APAs (two APAs tall) along the neutrino beam direction. A beam of high-energy neutrinos will be used for studies of neutrino oscillations with DUNE. For the full detector module, there are 150 APAs proposed in total, stacked two-tall and 25 along the beam direction. Because the spatial extent of the neutrino events under consideration is much smaller than the simulated workspace size, we expect the scaling down of the simulated geometry to have a negligible effect on the conclusions of this study.
| Channel | Expected Event Count |
| (all flavors) | 325.8 |
| 155.5 | |
| 67.3 | |
| 55.2 | |
| 47.8 | |
| 3300.0 |
@@ III.1 Neutrino-Electron Elastic Scattering
The most relevant interaction for directional information is when a neutrino elastically scatters off of an electron in the LAr. This interaction applies to all flavors but the largest cross section is for . A visualization of a Geant4 event is shown in Fig. 1 (b).
The direction of the scattered electron is highly correlated to the direction of the neutrino. In particular, the scattering angle is described by [vogel1989neutrino]:
where is the electron kinetic energy, is the neutrino energy, and is the electron mass. The distribution of is given by the differential cross section:
where is the Fermi coupling constant, and and are given according to neutrino flavor [vogel1989neutrino] in Tab. 2.
| Flavor | ||
@@
III.2 Electron-Neutrino Charged-Current Absorption Interactions